### v1

# Comparing next-generation detector configurations for high-redshift gravitational wave sources with neural posterior estimation

Abstract
The coming decade will be crucial for determining the final design and configuration of a global network of next-generation (XG) gravitational-wave (GW) detectors, including the Einstein Telescope (ET) and Cosmic Explorer (CE). In this study and for the first time, we assess the performance of various network configurations using neural posterior estimation (NPE) implemented in Dingo-IS–a method based on normalizing flows and importance sampling that enables fast and accurate inference. We focus on a specific science case involving short-duration, massive and high-redshift binary black hole (BBH) mergers with detector-frame chirp masses $\mathcal{M}_{\mathrm{d}}>100~\mathrm{M}_{\odot}$. These systems encompass early-Universe stellar and primordial black holes, as well as intermediate-mass black-hole binaries, for which XG observatories are expected to deliver major discoveries. Validation against standard Bayesian inference demonstrates that NPE robustly reproduces complex and disconnected posterior structures across all network configurations. For a network of two misaligned L-shaped ET detectors (2L MisA), the posterior distributions on luminosity distance can become multimodal and degenerate with the sky position, leading to less precise distance estimates compared to the triangular ET configuration. However, the number of sky-location multimodalities is substantially lower than the eight expected with the triangular ET, resulting in improved sky and volume localization. Adding CE to the network further reduces sky-position degeneracies, and the better performance of the 2L MisA configuration over the triangle remains evident.
[…]
show]filippo.santoliquido@gssi.it
I Introduction
Next-generation (XG) detectors such as the Einstein Telescope (ET; Punturo:2010zz; ET:2019dnz; Branchesi:2023mws) and Cosmic Explorer (CE; Reitze:2019iox; Evans:2023euw; Gupta:2023lga) will shape the future of gravitational-wave (GW) astronomy. Two of the greatest scientific breakthroughs expected with XG detectors are the observation of early-Universe binary black hole (BBH) mergers, potentially detectable up to redshifts of $z\lesssim 100$, and the detection of intermediate-mass black hole (IMBH) binary mergers, with source-frame total masses reaching up to $\lesssim 10^{4}\,\mathrm{M}_{\odot}$ (Hall:2019xmm). High-redshift BBHs include two main types of systems: (i) BBHs formed from Population III (Pop. III) stars, which have been extensively and recently studied as promising high-redshift GW sources of astrophysical origin (Hartwig:2016nde; Belczynski:2016ieo; kinugawa2020; liubromm2020; tanikawa2022; wang2022; Mestichelli:2024djn; Santoliquido:2024oqs). In particular, in santoliquido2023 we found that between $\sim 20$% and $\sim 70$% of detectable Pop. III BBHs with ET could merge at $z>8$; (ii) Primordial black holes (PBHs), originating from the collapse of large inhomogeneities during the radiation era (Zeldovich:1967lct; Hawking:1974rv; Chapline:1975ojl; Carr:1975qj), are also expected to form at high redshift and span a wide mass spectrum, including the stellar-mass regime accessible to XG observatories, and a broad range of merger rates (Ivanov:1994pa; Guth:2012we; Ivanov:1997ia; Blinnikov:2016bxu; DeLuca:2020qqa; Franciolini:2022tfm; Ng:2022agi; Carr:2020gox; Franciolini:2023opt).
In the $10^{2}$$10^{3}$ M range, IMBHs are thought to form primarily through stellar collisions (2002ApJ...576..899P; 2015MNRAS.454.3150G; 2016MNRAS.459.3432M; 2021A&A...652A..54A; 2023MNRAS.526..429A), hierarchical black hole mergers (Gerosa:2017kvu; Fishbach:2017dwv; Gerosa:2021mno; 2002MNRAS.330..232C; 2004ApJ...616..221G; 2019MNRAS.486.5008A; 2021A&A...652A..54A; Kritos:2022ggc), close interactions between stars and black holes (2017MNRAS.467.4180S; 2023MNRAS.521.2930R; 2023MNRAS.526..429A), or a combination of these mechanisms (2015MNRAS.454.3150G; 2023MNRAS.526..429A).
Accurate parameter estimation is essential for uncovering the origins of early-Universe BBHs and for determining the relative importance of the various formation pathways of IMBHs. However, the large expected detection rate of XG observatories, about $10^{5}$ events per year (Baibhav:2019gxm), poses unprecedented challenges for parameter estimation and data analysis (see also chapter 10 of Abac:2025saz).
To handle the vast amount of data expected, it is crucial to accelerate parameter estimation for individual sources, as current stochastic sampling methods, often requiring several hours per event (Smith:2019ucc), are inadequate for the era of XG detectors. Acceleration of parameter estimation can be pursued through three complementary strategies: simplifying the likelihood function (2010arXiv1007.4820C; Veitch:2014wba; Zackay:2018qdy; 2021PhRvD.104l3030L; Narola:2023men; Vinciguerra:2017ngf; Aubin:2020goo; 2021PhRvD.104d4062M; Allene:2025saz; 2015PhRvL.114g1104C; Smith:2021bqc; 2023PhRvD.107h4037T; Morras:2023pug; Smith:2021bqc; Narola:2024qdh; 2024arXiv241202651H), improving sampling efficiency (Williams:2021qyt; 2023MLS&T...4c5011W; nessai; 2025MNRAS.541..200P; 2023JOSS....8.5021W; 2023ApJ...958..129W; 2024PhRvD.110h3033W; Perret:2025wox; Negri:2025cyc; Prathaban:2025qgg), and employing methods based on deep neural networks (Cuoco:2020ogp).
In recent years, likelihood-free inference (or simulation-based inference)—particularly neural posterior estimation (NPE, Green:2020hst; Green:2020dnx; Dax:2021myb; Chatterjee:2022ggk; 2023mla..confE..34W; 2023PhRvL.130q1402L; Dax:2022pxd; DeSanti:2024oap; Lanchares:2025kpb; Marx:2025ioo; Srinivasan:2025etu; Kofler:2025dux)—has enabled much faster and accurate parameter estimation, producing posterior samples within minutes.
Dax:2021tsq implemented NPE in Dingo using conditional normalizing flows (Papamakarios:2019fms; Kobyzev_2021), which are invertible neural networks that transform a base distribution, typically a standard normal, into a more complex distribution which serves as an approximation of the posterior. Santoliquido:2025lot trained Dingo on the triangular configuration of ET and showed that NPE performs effectively in the XG-detector regime, delivering accurate parameter estimation with excellent computational efficiency.
The science goals of XG observatories critically depend on design choices that are effectively irreversible, including the detectors’ geometry, geographic location, and orientation (e.g., Branchesi:2023mws; Gupta:2023lga). Making the right decisions at the design stage is therefore crucial to maximizing the scientific potential of XG detectors.
This work presents the first quantitative assessment of how different XG detector configurations affect the parameter estimation performance for massive BBH mergers ($\mathcal{M}_{\mathrm{d}}>100~\mathrm{M}_{\odot}$), based on full Bayesian analyses. Leveraging NPE enables accurate evaluation of key metrics, including localization volumes and the presence of multimodal sky posteriors.
This manuscript is structured as follows. Section II.1 outlines the examined detector configurations, while Section II.2 introduces Dingo-IS, which combines NPE with importance sampling. Section II.3 defines the performance metrics used throughout the work. Section III.1 shows that Dingo-IS correctly recovers a large fraction of injections, between 85% and 96%, depending on the detector configuration. Section III.2 presents a single-injection study showing that a standard inference method and Dingo-IS yield statistically indistinguishable posteriors, with the latter requiring far less computational time. Section III.3 compares parameter-estimation performance across detector configurations, highlighting that two misaligned L-shaped ET detectors outperform the triangular ET for sky localization and volume. Finally, Section IV summarizes our conclusions.
II Methods
II.1 Detector and network configurations
We consider seven different XG detector configurations:
  • $\Delta$. One single triangular ET with 10-km arms located in Sardinia (Branchesi:2023mws).
  • 2L A. Two L-shaped ET detectors, each with 15-km arms, positioned in Sardinia and in the Meuse-Rhine Euro region (EMR), with parallel arms (Branchesi:2023mws).
  • 2L MisA. Same as 2L A, but with the detector arms misaligned by 45° (Branchesi:2023mws).
  • 2L MisA + LHI. Same as the 2L MisA configuration, but including three LIGO observatories—Livingston, Hanford, and India (Unnikrishnan:2013qwa)—operating at the mid-2030s A# LIGO sensitivity.
  • 1L + CE. An L-shaped ET detector with 15-km arms in Sardinia combined with a single CE detector with 40-km arms located near Hanford, WA (Reitze:2019iox). This configuration, introduced in Maggiore:2024cwf but not adopted as an official design by the ET Collaboration, is included here solely for comparison purposes.
  • $\Delta$ + CE. A single triangular ET detector located in Sardinia with 10-km arms observing with CE, as defined in 1L + CE.
  • 2L MisA + CE. Same as 1L + CE, but with a second L-shaped ET detector located at EMR, rotated by 45° relative to the first ET.
During the development of this work, the Lusatia region in Germany, near Kamenz (51.275°, 14.100°), was announced as the third official candidate site for ET. Its chord distance from the Sardinian site is approximately 1248 km, only about 82 km longer than the distance between the EMR region and Sardinia. Given this minimal difference, we consider the two sites effectively equivalent in hosting the second L-shaped interferometer, for the specific science case considered here.
Additional details, including coordinates, chord distances between interferometers and further information on the amplitude spectral densities (ASDs) are provided in Appendix A.
II.2 Dingo-IS
We conduct our analysis using NPE implemented in Dingo, where a conditional probabilistic neural network $q_{\phi}(\theta|d)$ with tunable parameters $\phi$ approximates the true posterior $p(\theta|d)$. Dingo is trained on simulated datasets $(\theta,d)$, where the parameters $\theta$ are sampled from the priors and the corresponding data $d$ consist of a GW signal embedded in stationary Gaussian noise, $d=h(\theta)+n$.
The prior distribution is the same as that adopted in Santoliquido:2025lot and listed in Appendix B, along with the choices of waveform parameters and approximants.
Compared to the network architecture in Santoliquido:2025lot, where all Dingo hyperparameters were kept at their default values, we enlarge the embedding network that compresses the input strain data into a set of features. Specifically, the output dimensionality is increased from 128 to 256 features, and several additional hidden layers are added, resulting in a total of $4\times 10^{8}$ learnable parameters.
Appendix C shows the learning curves for each considered detector configuration, along with the corresponding probability–probability plots.
After training, Dingo can rapidly generate approximate posterior samples. These samples, however, may deviate from the true posterior. A practical way to assess and refine them is via importance sampling (https://doi.org/10.1002/wics.56; Payne:2019wmy; Elvira_2021; Ashton:2025xba). Combining this technique with Dingo leads to Dingo-IS (Dax:2022pxd), which assigns weights to a set of $n$ samples $\theta_{i}$ drawn from the proposal distribution $q_{\phi}(\theta|d)$:
$w_{i}=\frac{\mathcal{L}(d|\theta_{i})\pi(\theta_{i})}{q_{\phi}(\theta_{i}|d)},$ (1)
where $\mathcal{L}(d|\theta)$ is the likelihood function defined in Appendix D. Ideally, all weights $w_{i}$ would be equal; in practice, the sample efficiency quantifies how well $q_{\phi}(\theta|d)$ approximates the target distribution:
$\epsilon=\frac{n_{\mathrm{eff}}}{n}\in\left[\frac{1}{n},\,1\right],$ (2)
with the effective number of samples given by $n_{\mathrm{eff}}=(\sum_{i}w_{i})^{2}/\sum_{i}w_{i}^{2}$ (Kong). The lower bound $\epsilon=1/n$ corresponds to the extreme case where one weight dominates and all others are negligible, which is the minimum possible as the weights are normalized to have mean one.
We compare the results obtained with Dingo-IS to those derived using standard inference methods implemented in Bilby (see Appendix D for additional details).
II.3 Metrics to assess detector performance
We define the metrics to evaluate the performance in estimating parameters of the various configurations of XG detectors.
The information gain ($I$) is defined as the Kullback–Leibler divergence (KL; 10.1214/aoms/1177729694) between the prior and the posterior (2022RNAAS...6...89B). In particular, we compute the KL divergence of the posterior from the prior:
$I=\int p(\theta|d)\log_{2}\frac{p(\theta|d)}{\pi(\theta)}\mathrm{d}\theta,$ (3)
which quantifies the extent to which the prior volume has been constrained by the data–by a factor $2^{I}$ (Skilling:2004pqw). As opposed to single- or few-parameter metrics, the information gain fully characterizes the size of the 11-dimensional posterior distribution, accounting for correlations across any parameters. Equation 3 is evaluated using a Monte Carlo integral. Additional details on this are provided in Appendix E.
Because Dingo-IS provides full parameter estimates for each source, we can define precise metrics to assess how different detector networks perform in source localization. For each event, we compute the area (in deg2) enclosed by the smallest 90% confidence region. The sky maps are sampled using equal-area HEALPix (Hierarchical Equal Area isoLatitude Pixelization; Gorski:1999rt; Gorski:2004by) pixels, where each of the $N$ pixels carries a posterior probability $p_{i}$ that the source lies within that pixel.
We employ the ligo-skymap library (Singer:2015ema; Singer:2016eax; Singer:2016erz), which also identifies the number of disconnected sky modes (or probability islands) within a given area using a flood-fill algorithm (foley1996computer), which finds all pixels connected to a starting one. Additionally, by including the posterior on luminosity distance, we compute the 90% highest probability density comoving volume ($\Delta V^{c}_{90\%}$, see Section 3 of Singer:2016eax).
The other reported interval widths, for example $\Delta x/x^{\mathrm{inj}}$, indicate the relative deviation of a parameter from its injected value. Here, $\Delta x$ corresponds to the square root of the associated covariance matrix element, representing the $1\sigma$ uncertainty in the estimate.
III Results
III.1 High signal-to-noise ratio vs. high sample efficiency
Figure 1: Top panel: Injections as function of sample efficiency ($\epsilon$, $x$-axis) and optimal SNR ($\rho$, $y$-axis) for each detector configuration (color-coded). Filled markers indicate injections with sample efficiencies $1\%$. Percentages in parentheses denote the fraction of sources with sample efficiency above $1\%$, while the horizontal colored lines indicate the median $\rho$ for sources exceeding this threshold. Bottom panel: Injections as function of detector-frame chirp mass ($x$-axis) and luminosity distance (left $y$-axis), with corresponding redshifts (right $y$-axis). Colored markers indicate events with sample efficiency greater than $1\%$ in all considered detector configurations. See Section III.1 for details.
We generate $1000$ sets of BBH merger parameters sampled from the prior and use the same samples for all configurations.
From the Dingo proposal distribution $q_{\phi}(\theta|d)$, we draw $10^{5}$ posterior samples and compute the corresponding importance weights (see Section II.2). For each injected signal, we then evaluate the sample efficiency. To ensure unbiased estimates of the inferred parameters, we require at least $10^{3}$ effective samples per source, corresponding to a sample efficiency >1>1 %. The top panel of Figure 1 shows that a large fraction of injected signals achieves a sample efficiency greater than $1\%$, ranging from $96\%$ for $\Delta$ configuration to a minimum of $85\%$ for 2L MisA + CE. These values remain consistently high, with even better performance for the $\Delta$ configuration than previously reported in Santoliquido:2025lot. This improvement arises from the increased dimensionality of the embedding network (see Section II.2), emphasizing the crucial role of this component within the NPE framework.
The top panel of Figure 1 also shows that $50\%$ of the events have an optimal signal-to-noise ratio (SNR, see Appendix D) exceeding $40$ for the $\Delta$ configuration and up to $70$ for 2L MisA + CE, while still achieving sample efficiencies above $1\%$. This demonstrates that the sample efficiency remains high even for high-SNR events, highlighting that NPE may represent a promising approach to address the parameter estimation challenges posed by XG detectors.
To ensure an unbiased evaluation of parameter estimation performance, we include only injections with sample efficiency exceeding $1\%$ across all seven detector configurations. The bottom panel of Figure 1 illustrates the parameter space of the retained sources, which account for 78% of the total, showing that events with $\mathcal{M}_{d}\lesssim 150\,\mathrm{M}_{\odot}$ and $D_{\mathrm{L}}\lesssim 6$ Gpc are almost entirely excluded. These sources have low sample efficiencies because their high SNR makes them more challenging for NPE to model, as their posteriors are much narrower than the prior. This behavior is further evident in the top panel of Figure 1, where sample efficiency drops sharply for sources with optimal SNR around $\sim 1000$.
We are currently exploring several strategies to improve sample efficiency in regions of parameter space where performance remains below threshold, including training Dingo-IS with neural population priors (e.g., Wouters:2025csq). The results of these efforts will be presented in future work.
We further restrict the dataset to sources with optimal SNR $\rho>8$.
III.2 Multimodal posterior of a single event
Figure 2: Top panel: Marginalized one-dimensional posterior distributions for the luminosity distance, recovered with Dingo-IS for an event observed with the $\Delta$ (dark blue) and 2L MisA (light blue) configurations. The results are compared with Bilby (orange). The vertical line marks the injected luminosity distance. The remaining parameters are shown in Appendix D. Middle panel: Posterior samples obtained with Dingo-IS in the $\Delta$ configuration for right ascension (ra) and declination (dec), color-coded by luminosity distance. The red cross marks the injected sky position. We also report the sky-localization area ($\Delta\Omega_{90\%}$) and comoving-volume localization ($\Delta V^{c}_{90\%}$) errors. Bottom panel: Same as the middle panel, but for the 2L MisA configuration. See Section III.2 for details.
The top panel of Figure 2 shows that Dingo-IS yields posterior distributions that are statistically indistinguishable from those obtained using Bilby (see Appendix D for a formal validation). Crucially, this accuracy is achieved at a dramatically lower computational cost: Dingo-IS requires only a few minutes per source, whereas Bilby needs more than 22 hours (50 hours) to converge for this event with the $\Delta$ (2L MisA) configuration.
Figure 2 also shows that Dingo-IS accurately captures the complex structure of multimodal posteriors. These multimodalities include both the eight-fold sky degeneracy arising from the triangular geometry of ET, extensively discussed in Santoliquido:2025lot, and modes introduced by the 2L MisA configuration, affecting both sky localization and luminosity distance.
The bottom panel of Figure 2 shows that for the 2L MisA configuration the sky location is correlated with luminosity distance: distinct islands of probability on the sky correspond to different modes in $D_{\mathrm{L}}$. For this source, we identify three such modes. The occurrence rate of multimodal posteriors in luminosity distance is discussed in Appendix F and shown in Figure 10. The 2L MisA (2L A) configurations produce a multimodal $D_{\mathrm{L}}$ posterior more frequently than the $\Delta$ configuration, with rates of about 20% compared to $\sim$2%, respectively.
These multimodalities are expected when observing short-duration, high-mass BBHs with both the 2L A and 2L MisA configurations of ET. They arise from the combined effect of the short baseline between the two interferometers ($\sim$1000 km) and the low merger frequencies of massive BBHs (typically $\lesssim 50$ Hz). We provide an illustrative example in Appendix G, showing that for a signal with higher merger frequencies, the multimodalities in both sky position and distance disappear.
Despite the multimodality in $D_{\mathrm{L}}$, the number of sky-position modes is significantly reduced compared to the triangular configuration, resulting in improved sky-area and comoving-volume localization for the 2L MisA configuration.
Additional details, including the same source as observed with the 2L MisA + CE configuration, are provided in Appendix D.
III.3 Parameter estimation performance
Figure 3: Cumulative distributions of events as a function of the information gain (top-left panel, see also Section II.3), the relative variations in luminosity distance (top-right panel, $\Delta D_{\mathrm{L}}/D^{\mathrm{inj}}_{\mathrm{L}}$), sky (bottom-left panel, $\Delta\Omega_{90\%}$, see also Table 1), and comoving volume localization (bottom-right panel, $V^{c}_{90\%}$) for all considered XG detector configurations (color-coded). See Section III.3 for details and Appendix H for the remaining parameters.
The top-left panel of Figure 3 shows the parameter estimation performance of the considered detector configurations, measured by the information gain, revealing a clear ranking. Among the ET configurations ($\Delta$, 2L A, 2L MisA), when no other XG detectors are observing simultaneously, the 2L MisA setup performs best. Moreover, adding three LIGO observatories operating at A# sensitivity to the 2L MisA configuration further increases the information gain, underscoring the capability of this global network to estimate the parameters of the specific sources examined in this study—namely, both massive and high-redshift BBHs. The information gain further increases when 1L ET observes jointly with CE, as the longer baseline enhances both sky and volume localization. However, the separation between the 2L MisA detectors also plays a role: when observing together with CE, the 2L MisA + CE configuration outperforms $\Delta$ + CE.
The top-right panel of Figure 3 shows that the luminosity distance is best measured with the $\Delta$ configuration, outperforming both the 2L A and 2L MisA configurations. Compared to Branchesi:2023mws (see their Figure 5), we find a reversed hierarchy in detector performance. We stress that our analysis is restricted to both massive and high-redshift BBH mergers. For this class of sources, the 2L configurations give rise to multimodal posteriors in $D_{\mathrm{L}}$ (see Figure 2 and Appendix F), an effect that is not captured with Fisher-matrix–based analyses.
Nevertheless, the 2L MisA configuration leads to better sky and volume localization, allowing it to outperform the $\Delta$ and 2L A setup in these metrics (bottom panels of Figure 3). Improving the accuracy of the localization volume critically enhances the prospects for dark siren cosmology, which relies on the statistical association between the GW source and galaxies contained within the reconstructed localization volume (Schutz:1986gp; DelPozzo:2011vcw; Chen:2017rfc; Libanore:2020fim; Gair:2022zsa; Bosi:2023amu; Borghi:2023opd).
In Appendix H, we present the distribution of optimal SNR for individual sources, along with the performance of the detector configurations on other parameters—including the detector-frame chirp mass, mass ratio, aligned spins of both components, and inclination angle—which lead to broadly similar conclusions.
$\Delta\Omega_{90\%}$ <10<10 deg2 <100<100 deg2 <1000<1000 deg2
$\Delta$ 0.7% 8.5% 26.2%
2L A 0.3% 4.4% 24.3%
2L MisA 3.6% 15.0% 34.8%
2L MisA + LHI 9.5% 29.6% 50.9%
1L + CE 17.3% 42.2% 69.7%
$\Delta$ + CE 31.0% 68.0% 93.2%
2L MisA + CE 36.7% 71.5% 94.2%
Table 1: Percentage of injections (color-coded) with estimated sky localization errors within the specified areas (columns) for different detector configurations (rows). See Section III.3 for details.
Table 1 summarizes the sky localization performance for all detector configurations, showing a clear improvement from the 2L A configuration, which performs worst, to 2L MisA + CE, which achieves the best results. In the latter case, more than 70% of sources are localized within 100 deg2.
Figure 4: Percentage of sky maps ($y$-axis) with $\Delta\Omega_{90\%}\leq 100$ deg2 exhibiting one to eight or more disconnected modes (color-coded), for all detector configurations ($x$-axis). See Section III.3 for details.
Figure 4 shows that more than 80% of events exhibit eight sky modes in the $\Delta$ configuration. This fraction is substantially reduced for the two L-shaped detector configurations: in the 2L MisA setup, the occurrence of eight sky modes drops to about 20% of the events.
The 2L MisA configuration exhibits an eight-sky-mode pattern for events lying close to the local horizon corresponding to the midpoint between the two L-shaped interferometers. For such events, the time-of-arrival difference is negligible, effectively making the two interferometers colocated (Marsat:2020rtl).
The 2L A setup localizes nearly 30% of the sources into two distinct sky modes, outperforming the 2L MisA configuration, where this fraction remains below 10%. However, the average sky-localization area per mode is significantly larger in 2L A compared to 2L MisA.
None of the ET configurations are able to localize events into a single sky mode unless ET operates as part of a global network that includes either the three LIGO interferometers at A# sensitivity (including LIGO–India) or CE.
The performance of ET configurations in estimating parameters of compact binary coalescences arises from the interplay between polarization measurements and a baseline that enables triangulation (Fairhurst:2009tc). Breaking the degeneracy between $D_{\mathrm{L}}$ and $\theta_{\mathrm{jn}}$ requires measuring both polarizations (Usman:2018imj), and the ability to do so is quantified by the alignment factor (Klimenko:2006rh; deSouza:2023gjv; Mascioli:2025cnw). A network composed of two L-shaped interferometers that are perfectly aligned provides no information on the cross-polarization. In contrast, the $\Delta$ configuration exhibits the highest alignment factor, yielding the most accurate polarization measurements. However, the sky maps produced by this configuration feature eight distinct sky modes owing to the colocated nature of the detectors (Singh:2020wsy; Singh:2021bwn; Santoliquido:2025lot). As a result, the 2L MisA configuration offers a trade-off, enabling polarization measurements while reducing the number of sky modes (see Figure 4).
IV Conclusions
Given the unique sensitivity that ET and other XG gravitational wave detectors will provide to early-Universe stellar and primordial black holes, as well as to intermediate-mass black-hole binaries, we focus our analysis on BBH mergers with detector-frame chirp masses $\mathcal{M}_{\mathrm{d}}>100~\mathrm{M}_{\odot}$. Although this study addresses a more restricted science case than the full range of sources observable with XG detectors (Branchesi:2023mws; Abac:2025saz), it targets a regime with strong potential for breakthrough discoveries.
We assess the performance of seven configurations of XG detectors using neural posterior estimation enhanced with importance sampling, as implemented in Dingo-IS (Dax:2022pxd).
We conduct a large injection campaign by sampling 1000 BBH merger parameters from the prior. As shown in Figure 1, Dingo-IS achieves excellent performance, recovering 85–96% of signals with a sample efficiency above 1%, depending on the detector configuration and for median SNRs between 40 and 70. This demonstrates that its efficiency remains high even for high-SNR events.
In Section III.2, we present a single-injection study showing that Dingo-IS accurately reproduces complex, disconnected posteriors consistent with standard stochastic-sampling methods, but at a fraction of the computational cost—few minutes instead of several hours.
Our analysis shows that a network of two misaligned L-shaped ET detectors (2L MisA) can produce multimodal luminosity–distance posteriors (see top panel of Figure 2), and its distance estimates are generally less precise than those obtained with the triangular ET design ($\Delta$). However, Figure 3 demonstrates that the 2L MisA outperforms the $\Delta$ configuration in terms of information gain, sky localization, and volume reconstruction. The $\Delta$ configuration’s larger number of sky modes leads to poorer sky-localization performance, which in turn dominates its volume uncertainty.
For high-mass and high-redshift BBH events, the fraction of sources recovered within a single sky mode remains low for all ET configurations ($\Delta$, 2L A, and 2L MisA). It surpasses 50% when CE is added to the global XG network and exceeds 70% when, instead, three LIGO detectors operate at A# sensitivity, as shown in Figure 4.
We thank Stephen R. Green, Francesco Iacovelli, Luca Reali, Alexandre Toubiana, Manuel Arca Sedda, Michele Maggiore, Archisman Gosh, Elisa Maggio and Emanuele Berti for useful discussions. F.S. has been funded by the European Union – NextGenerationEU under the Italian Ministry of University and Research (MUR) ”Decreto per l’assunzione di ricercatori internazionali post-dottorato PNRR” - Missione 4 ”Istruzione e Ricerca” Componente 2 ”Dalla Ricerca all’Impresa” del PNRR - Investimento 1.2 “Finanziamento di progetti presentati da giovani ricercatori” - CUP D13C25000700001. M.B. and J.H. acknowledge support from the Astrophysics Center for Multi-messenger Studies in Europe (ACME), funded under the European Union’s Horizon Europe Research and Innovation Program, Grant Agreement No. 101131928. The data analysis has been performed on the HPC workstations ”Merac2A” financially supported by the MERAC Foundation and the European Astronomical Society through the 2023 MERAC Prize and the MERAC Support to research (PI Manuel Arca Sedda). The research leading to these results has been conceived and developed within the Einstein Telescope Observational Science Board ET-0611A-25.
Appendix A Detector and network configurations
We provide further details on the detector and network configurations considered in this study. We limit the frequency range from $f_{\mathrm{min}}=6$ Hz to $f_{\mathrm{max}}=256$ Hz, using 8-second time segments, which results in frequency bins of size $\mathrm{d}f=0.125$ Hz. Setting the lower frequency limit to $f_{\mathrm{min}}=6$ Hz restricts the duration of all processed signals to less than 8 seconds. This approximation slightly underestimates ET performance, given its expected sensitivity down to $\sim 2$ Hz, but remains adequate for CE, whose sensitivity declines markedly below 5 Hz.
All ET configurations ($\Delta$, 2L A, and 2L MisA) adopt the HFLF cryogenic ASD curve (et), which incorporates both a high-frequency (HF) instrument and a cryogenic low-frequency (LF) instrument. Locations and orientations for the ET configurations are taken from Branchesi:2023mws. The ASDs for the three LIGO observatories (Livingston, Hanford, and India) at A# sensitivity is taken from ligo-india-asd, while the coordinates and orientation of the Indian interferometer from ligo-india. The location of CE is assumed to be the near the LIGO-Hanford Observatory (LHO), while the ASD for the 40-km arms in the baseline configuration is taken from ce_asd. Figure 5 shows the amplitude spectral densities (ASDs) used in this study for each detector over the specified frequency range.
Table 2 summarizes the locations, orientations, and chord distances between interferometers of the various detector configurations. Chord distances are calculated assuming the Earth is a perfect sphere with radius $R_{\oplus}=6378.1366$ km.
Figure 5: Amplitude spectral densities (ASDs) adopted in this work for different XG detectors. See Section II.1 and Appendix A for details.
Detector Location $x$-arm azimuth Chord distances km
$\Delta$ ET-1: 40.517°, 9.417° ET-1: 250.567° ET-1 $\leftrightarrow$ ET-2 10
ET-2: 40.432°, 9.377° ET-2: 130.567° ET-2 $\leftrightarrow$ ET-3 10
ET-3: 40.448°, 9.493° ET-3: 10.567° ET-3 $\leftrightarrow$ ET-1 10
2L A ET-S: 0° ET-S $\leftrightarrow$ ET-EMR 1166
ET-S: 40.517°, 9.417° ET-EMR: 0°
2L MisA ET-EMR: 50.723°, 5.921° ET-S: 0°
ET-EMR: 45.0°
Livingston (LLO) 30.563°, 90.774-90.774° 197.7165° ET-S $\leftrightarrow$ LLO 7994
ET-EMR $\leftrightarrow$ LLO 7382
Hanford (LHO) 46.455°, 119.408-119.408° 125.9994° LLO $\leftrightarrow$ LHO 3002
ET-S $\leftrightarrow$ LHO 8352
ET-EMR $\leftrightarrow$ LHO 7498
India (LIO) 19.613°, 77.031° 117.6157° LIO $\leftrightarrow$ ET-S 6436
LIO $\leftrightarrow$ ET-EMR 6671
LIO $\leftrightarrow$ LLO 11488
LIO $\leftrightarrow$ LHO 10593
CE 46.455°, 118.592-118.592° 154.4305° CE $\leftrightarrow$ ET-EMR 7470
CE $\leftrightarrow$ ET-S 8323
CE $\leftrightarrow$ ET-1
Table 2: Locations (latitude and longitude), orientations ($x$-arm azimuth), and chord distances for the different configurations considered in this work. See Section II.1 and Appendix A for details. Notes: The $x$-arm azimuth is measured counterclockwise from East. For all detectors, the $y$-arm azimuth is given by $x$-arm azimuth + 90°, except for the $\Delta$ configuration, where the $y$-arm azimuth = $x$-arm azimuth + 60°.
Appendix B Waveforms and priors
We define spin-aligned, quasi-circular waveforms defined by the parameter set $\theta=\{\mathcal{M}_{\mathrm{d}}$, $q$, $D_{\mathrm{L}}$, $\mathrm{dec}$, $\mathrm{ra}$, $\theta_{\mathrm{jn}}$, $\psi$, $t_{\mathrm{geocent}}$, $\chi_{1}$, $\chi_{2}$, $\phi_{c}\}$, where $\mathcal{M}_{\mathrm{d}}$, $q$, and $D_{\mathrm{L}}$ denote the detector-frame chirp mass, mass ratio, and luminosity distance, respectively; $\mathrm{ra}$ and $\mathrm{dec}$ represent the right ascension and declination of the source; $\theta_{\mathrm{jn}}$ denotes the angle between the line of sight and the total angular momentum $\vec{J}$. In the absence of precession, as assumed in this work, $\vec{J}$ is equal to the orbital angular momentum of the binary $\vec{L}$, and thus $\theta_{\mathrm{jn}}$ corresponds to the inclination angle $\iota$; $\psi$ is the polarization angle; $t_{\mathrm{geocent}}$ is the coalescence time at the Earth’s center; and $\phi_{c}$ is the phase at coalescence. The waveforms also depend on the aligned-spin parameters $\chi_{1}$ and $\chi_{2}$, defined as $\chi_{i}=a_{i}\cos\theta_{i}$, where $a_{i}=|\chi_{i}|$ is the spin magnitude and $\theta_{i}$ is the angle with respect to the orbital angular momentum.
We employ frequency-domain waveforms with a reference frequency fixed at $f_{\mathrm{ref}}=20$ Hz (Varma:2021csh), using the IMRPhenomXPHM waveform approximant (PhysRevD.103.104056), which includes the subdominant modes $(\ell,|m|)=(2,1),(3,3),(3,2),(4,4)]$ in addition to the dominant $(\ell,|m|)=(2,2)]$ mode.
The prior distributions $\pi(\theta)$ employed in training Dingo are reported in Table 3. In particular, for the detector-frame chirp mass and mass ratio, we choose a prior uniform in primary and secondary masses (GWFish_Priors_Tutorial), therefore:
$\pi(\mathcal{M}_{\mathrm{d}})\propto\mathcal{M}_{\mathrm{d}},$ (B1)
where $\mathcal{M}_{\mathrm{d}}\in[40,~1100]$ M; and:
$\pi(q)\propto\frac{(1+q)^{2/5}}{q^{6/5}},$ (B2)
where $q\in[0.125,~1]$. For the luminosity distance, we choose a prior uniform in comoving volume and source-frame time,
$\pi(D_{\mathrm{L}})\propto\frac{\mathrm{d}V_{c}}{\mathrm{d}D_{\mathrm{L}}}\frac{1}{1+z},$ (B3)
where $D_{\mathrm{L}}\in[5,~500]$ Gpc, assuming cosmological parameters as provided by Planck:2018vyg. For the aligned spin component $\chi_{i}$, we form a joint prior as follows:
$\pi(\chi_{i})=\int_{0}^{0.9}\mathrm{d}a_{i}\int_{-1}^{1}\mathrm{d}\cos\theta_{i}\pi(a_{i})\pi(\cos\theta_{i})\delta(\chi_{i}-a_{i}\cos\theta_{i}),$ (B4)
where $\pi(a_{i})=\mathcal{U}(0,0.9)$ and $\cos\theta_{i}=\mathcal{U}(-1,1)$ (lange2018rapidaccurateparameterinference; Ashton:2018jfp).
By convention, we set the Earth’s orientation and the positions and orientations of interferometers to those corresponding to the reference time $t_{\mathrm{ref}}=1126259462.391$ s, which coincides with the GPS trigger time of GW150914 (Romero-Shaw:2020owr). The merger time ($t_{\mathrm{geocent}}$) of each event is randomly sampled around this reference time. Fixing $t_{\mathrm{ref}}$ does not entail any loss of generality for short signals, as any time offset can be recovered.
Parameter Units Prior
$\mathcal{M}_{\mathrm{d}}$ M Eq. B1
$q$ Eq. B2
$D_{\mathrm{L}}$ Gpc Eq. B3
ra rad $\mathcal{U}(0,2\pi)$
$\cos$ dec $\mathcal{U}(-1,1)$
$\sin\theta_{\mathrm{jn}}$ $\mathcal{U}(0,1)$
$\phi_{c}$ rad $\mathcal{U}(0,2\pi)$
$\psi$ rad $\mathcal{U}(0,\pi)$
$t_{\mathrm{geocent}}$ s $\mathcal{U}(-0.1,+0.1)$
$\chi_{i}$ Eq. B4
Table 3: Parameters defining GW signals and their adopted priors. $\mathcal{U}(a,b)$ represents a uniform distribution between $a$ and $b$. See Appendix B for details.
Appendix C Learning curves and p-p plots
Figure 6 shows that training and test losses remain consistent across all considered detector configurations, indicating the absence of overfitting. We observe an improvement in the training process compared to Figure 3 of Santoliquido:2025lot, where the test losses exhibited significant noise before epoch $\sim 200$.
Figure 7 displays the probability–probability (P–P) plots for the different detector configurations. P–P plots are commonly used to assess the reliability of inference algorithms (Cook01092006; Veitch:2014wba; talts2020validatingbayesianinferencealgorithms; 2020PhRvD.102j4057G), by verifying whether the fraction of injected parameters recovered within a given credible interval follows a uniform distribution. This method also enables the computation of $p$-values for each parameter through a Kolmogorov–Smirnov test. The results reported in Table 4 show good overall consistency. The plots in Figure 7 demonstrate that the improved training procedure produces an unbiased proposal distribution $q_{\phi}(\theta|d)$, as evidenced by the curves lying close to a uniform distribution.
Figure 6: The log loss as a function of training epochs shown for all detector configurations considered in this study, with the solid line representing the training set and the dashed line representing the test set. See Appendix C for further details.
[…]
$\Delta$ 2L A 2L MisA 2L MisA + LHI 1L + CE $\Delta$ + CE 2L MisA + CE
$\mathcal{M}_{\mathrm{d}}$ 0.72 0.12 0.92 0.29 0.02 0.01 0.91
$q$ 0.35 0.03 0.59 0.23 0.04 0.47 0.76
$D_{\mathrm{L}}$ 0.08 0.12 0.42 0.03 0.04 0.84 0.47
ra 0.79 0.53 0.73 0.59 0.71 0.90 0.57
dec 0.80 0.71 0.76 0.03 0.19 0.47 0.34
$\theta_{\mathrm{jn}}$ 0.92 0.96 0.09 0.37 0.34 0.72 0.98
$\psi$ 0.79 0.72 0.92 0.23 0.25 0.43 0.35
$t_{\mathrm{geocent}}$ 0.55 0.30 0.12 0.27 0.37 0.29 0.08
$\chi_{1}$ 0.55 1.00 0.21 0.41 0.19 0.61 0.83
$\chi_{2}$ 0.21 0.28 0.47 0.06 0.42 0.15 0.43
$\phi_{c}$ 0.89 0.07 0.50 0.33 0.50 0.44 0.04
Table 4: $p$-values (color-coded) for each GW parameter (rows) across different detector configurations (columns). See Figure 7 and Appendix C for details.
Appendix D Dingo-IS vs. Bilby
2L MisA 2L MisA + CE
Parameters $\theta^{\mathrm{inj}}$ $\langle\mathrm{JSD}\rangle$ JSDthr $\theta$ $\langle\mathrm{JSD}\rangle$ JSDthr $\theta$
$\mathcal{M}_{\mathrm{d}}/$M $749$ 1.8 4.7 ${757}_{-7}^{+7}$ 2.9 5.5 ${751}_{-3}^{+3}$
$q$ $0.73$ 1.0 5.0 ${0.75}_{-0.02}^{+0.02}$ 2.0 4.7 ${0.73}_{-0.01}^{+0.01}$
$D_{\mathrm{L}}$/Gpc $13.8$ 1.7 4.9 ${18}_{-5}^{+1}$ 2.2 4 ${13.7}_{-0.3}^{+0.3}$
ra/rad $4.60$ 1.7 7.3 ${2.40}_{-0.15}^{+2.25}$ 0.0 0.0 ${4.60}_{-0.01}^{+0.01}$
dec/rad 0.40-0.40 1.8 5.6 ${0.16}_{-0.56}^{+0.12}$ 0.0 0.1 ${-0.40}_{-0.01}^{+0.01}$
$\theta_{\mathrm{jn}}$}/rad $1.06$ 0.4 2.0 ${2.04}_{-0.98}^{+0.05}$ 0.1 0.8 ${1.07}_{-0.01}^{+0.01}$
$\phi_{\mathrm{c}}$}/rad $0.09$ 0.5 2.1 ${0.11}_{-0.10}^{+6.17}$ 7.6 3.3 ${0.08}_{-0.06}^{+0.08}$
$\psi$}/rad $2.14$ 2.3 7.8 ${1.62}_{-0.10}^{+0.54}$ 0.0 0.7 ${2.13}_{-0.01}^{+0.01}$
$t_{\mathrm{geocent}}$/s $0.056$ 1.6 5.2 ${0.082}_{-0.029}^{+0.003}$ 2.9 6.6 ${0.056}_{-0.001}^{+0.001}$
$\chi_{1}$} $0.35$ 0.7 2.6 ${0.36}_{-0.07}^{+0.07}$ 2.3 3.9 ${0.36}_{-0.04}^{+0.04}$
$\chi_{2}$} $0.20$ 1.1 3.6 ${0.22}_{-0.10}^{+0.08}$ 2.4 6.6 ${0.21}_{-0.06}^{+0.06}$
$\rho$ 232 331
$m_{1,\mathrm{s}}$/M 364.2 ${317}_{-8}^{+66}$ ${366}_{-3}^{+3}$
$m_{2,\mathrm{s}}$/M 265.2 ${236}_{-6}^{+49}$ ${267}_{-3}^{+3}$
$z$ 1.8 ${2.18}_{-0.54}^{+0.09}$ ${1.77}_{-0.03}^{+0.03}$
${\Delta\Omega}_{90\%}$/deg2 235 1
${\Delta V}^{c}_{90\%}$/Mpc3 $4.1\times 10^{8}$ $7.2\times 10^{5}$
Sample efficiency/% 16 10
Dingo-IS $\log\mathcal{Z}(d)\pm\sigma$ $4046.006\pm 0.007$ $6180.05\pm 0.01$
Bilby $\log\mathcal{Z}(d)\pm\sigma$ $4046.29\pm 0.07$ $6179.26\pm 0.07$
Computing time/hours 51.5 8.8
Table 5: Injected parameters $\theta^{\mathrm{inj}}$ of the event shown in Figure 2, along with the median Jensen–Shannon divergence ($\langle\mathrm{JSD}\rangle$, in units of $10^{-4}$ nat), which quantifies the deviation between Dingo-IS and Bilby for one-dimensional marginal posteriors. For reference, the corresponding JSD threshold (JSDthr, in units of $10^{-4}$ nat) is also shown. The median and 90% credible interval of the recovered parameters $\theta$ with Dingo-IS are reported for two different detector configurations: 2L MisA and 2L MisA + CE. The bottom rows display the optimal SNR ($\rho$, Equation D4), source-frame primary ($m_{1,\mathrm{s}}$) and secondary ($m_{2,\mathrm{s}}$) masses, redshift ($z$), sky and comoving volume localization error ($\Delta\Omega_{90\%}$ and $\Delta V^{c}_{90\%}$), sample efficiency (Equation 2), log evidence and its uncertainty computed with Dingo-IS (Equation D5) and Bilby, and computing time in Bilby obtained using npool = 8 parallel processes. See Appendix D for details.
Figure 8: Marginalized one- and two-dimensional posterior distributions for all parameters for the event shown in Figure 2 observed with the 2L MisA configuration. Results from Dingo-IS (light blue) are compared with Bilby (orange). Vertical and horizontal lines indicate the true injected values (see Table 5) and contours represent the 68% and 95% credible regions. See Appendix D for further details.
Figure 9: Marginalized one- and two-dimensional posterior distributions for all parameters of the same event shown in Figure 8, this time observed with the 2L MisA + CE configuration. Results from Dingo-IS (light blue) are compared with Bilby (orange). Vertical and horizontal lines indicate the true injected values (see Table 5) and contours represent the 68% and 95% credible regions. See Appendix D for further details.
The parameters $\theta$ of a GW signal are inferred from strain data $d$ through Bayes’ theorem (vanderSluys:2007st; vanderSluys:2009bf; Thrane:2018qnx; Christensen:2022bxb):
$p(\theta|d)=\frac{\mathcal{L}(d|\theta)\pi(\theta)}{\mathcal{Z}(d)},$ (D1)
where $\mathcal{L}(d|\theta)$ is the likelihood function, $\pi(\theta)$ is the prior distribution (see Appendix B) and $\mathcal{Z}(d)=\int\mathrm{d}\theta\mathcal{L}(d|\theta)\pi(\theta)$ is the evidence.
Assuming stationary Gaussian noise, the likelihood in Equation D1 takes the form (Finn:1992wt; Veitch:2014wba; Romano:2016dpx; Ashton:2018jfp):
$\mathcal{L}(d|\theta)\propto\exp\left[\sum_{i}\left((d_{i}|h_{i}(\theta))-\frac{1}{2}(h_{i}(\theta)\,|\,h_{i}(\theta))\right)\right],$ (D2)
where the index $i$ labels the detectors in the network, $h_{i}(\theta)$ denotes the GW signal, and $(\cdot|\cdot)$ the noise-weighted inner product, defined as
$(a|b)=4\,\mathrm{Re}\int_{f_{\mathrm{min}}}^{f_{\mathrm{max}}}\mathrm{d}f\,\frac{a^{*}(f)\,b(f)}{S(f)},$ (D3)
with $\mathrm{Re}$ indicating the real part and the complex conjugate. The quantity $\sqrt{S(f)}$ represents the noise ASD (see Appendix A for details). Equation D3 is also used to compute the optimal signal-to-noise ratio (SNR):
$\rho=\sqrt{(h(\theta^{\mathrm{inj}})|h(\theta^{\mathrm{inj}}))},$ (D4)
where $h(\theta^{\mathrm{inj}})$ denotes the GW signal evaluated with injected parameters. For a network of detectors, the total optimal SNR is obtained by adding the individual detector SNRs in quadrature.
To sample from Equation D1, we employ Bilby (Romero-Shaw:2020owr), a Python library designed for Bayesian inference with the nested sampling algorithm nessai (nessai). Thanks to its use of normalizing flows, nessai achieves rapid convergence. We fix the number of live points to n_live = 10_000.
For the source shown in Figure 2 and observed with the 2L MisA configuration, the default nessai settings yield suboptimal performance. To improve robustness, we increase the volume_fraction from 0.95 to 0.98. This parameter controls the fraction of the total probability enclosed by the contour in the latent space—the base space from which samples are initially drawn before being mapped to physical space through the normalizing flow (Williams:2021qyt). Although a higher volume fraction improves performance, it also increases computational cost, which explains the long runtime ($\sim$ 50 hours) required for this source (see Table 5).
To quantify the agreement between posterior samples obtained with Dingo-IS and Bilby, we employ the Jensen–Shannon divergence (JSD; 61115), a symmetrized variant of the Kullback–Leibler divergence (10.1214/aoms/1177729694). The JSD provides a measure of the similarity between two probability distributions, taking values from 0 nat, when the distributions are identical, up to $\ln(2)=0.69$ nat for completely distinct distributions.
To determine a threshold for the JSD beyond which one-dimensional posteriors can be regarded as statistically different, we follow the procedure outlined in Santoliquido:2025lot. Using Dingo-IS, we generate 100 independent sets of $10^{5}$ posterior samples. For each parameter $\theta$, we compute the JSD across all 4950 unique pairs of these sets. Because all samples originate from the same underlying distribution, the resulting JSD values capture statistical fluctuations rather than meaningful discrepancies. We adopt the 95th percentile of these values as the threshold for each parameter (JSDthr in Table 5). Finally, we compare the one-dimensional posteriors obtained with Bilby to each Dingo-IS set, and verify whether the median JSD lies below the corresponding threshold ($\langle\mathrm{JSD}\rangle$ in Table 5).
Table 5 lists the JSD values computed for the marginal one-dimensional posteriors of the source analyzed in Figure 2. The results show that the posteriors obtained with Bilby and Dingo-IS are statistically indistinguishable for both the 2L MisA and 2L MisA + CE configurations. This excellent agreement is further illustrated in Figures 8 and 9, which display the marginalized one- and two-dimensional posteriors for all parameters for the 2L MisA and 2L MisA + CE networks, respectively.
Table 5 also reports the Bayesian evidence $\mathcal{Z}(d)$ defined in Equation D1, along with its uncertainty estimated by averaging the weights (Equation 1) obtained with Dingo-IS (mcbook):
$\mathcal{Z}(d)\pm\sigma=\frac{1}{n}\sum_{i}w_{i}\left(1\pm\sqrt{\frac{1-\epsilon}{n-\epsilon}}\right),$ (D5)
where $n$ denotes the total number of samples drawn from the proposal distribution $q_{\phi}(\theta|d)$, and $\epsilon$ is the sampling efficiency defined in Equation 2. The log evidences computed with Dingo-IS and Bilby are consistent within statistical uncertainties.
Appendix E Information gain
We compute Equation 3 through a Monte Carlo integration:
$I=\int p(\theta|d)\log_{2}\frac{p(\theta|d)}{\pi(\theta)}\mathrm{d}\theta\approx\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\log_{2}\frac{p(\theta_{i}|d)}{\pi(\theta_{i})},$ (E1)
where $N_{s}=3\times 10^{4}$ denotes the number of samples drawn from the posterior. In our case, these samples are obtained through sampling–importance–resampling (SIR; Rubin1988UsingTS) based on the importance weights defined in Equation 1. Using Bayes’ theorem, we can rewrite the ratio inside the logarithm as:
$I=\frac{1}{N_{s}}\sum_{i}^{N_{s}}\log_{2}\frac{p(\theta_{i}|d)}{\pi(\theta_{i})}=\frac{1}{N_{s}}\sum_{i}^{N_{s}}\log_{2}\frac{\mathcal{L}(d|\theta_{i})\cancel{\pi(\theta)}}{\mathcal{Z}(d)\cancel{\pi(\theta)}}=-\log_{2}\mathcal{Z}(d)+\frac{1}{N_{s}}\sum_{i}^{N_{s}}\log_{2}\mathcal{L}(d|\theta_{i}),$ (E2)
where the log evidence $\log_{2}\mathcal{Z}(d)$ is evaluated through the importance sampling weights (Equation D5). We estimate the variance associated with the Monte Carlo estimator of the information gain (2003itil.book.....M; Talbot:2023pex; Heinzel:2025ogf) as follows:
$\sigma_{I}^{2}=\frac{1}{N_{s}}\left[\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\left(-\log_{2}\mathcal{Z}(d)+\log_{2}\mathcal{L}(d|\theta_{i})\right)^{2}-I^{2}\right],$ (E3)
where $I$ is defined in Equation E2 and $\sigma_{I}^{2}<10^{-3}$ bits for all detector configurations.
Appendix F Multimodal luminosity distance
Figure 10: Percentage of luminosity distance posteriors ($y$-axis) that are multimodal (dark blue) or unimodal (light blue) for all detector configurations ($x$-axis). See Appendix F for details.
We apply Hartigan’s dip test for unimodality (fb099965-2cf5-3955-b5e5-077ac9290f24; 4c99dc66-50c0-3bf4-8c15-3b2f3d6ec1eb) to the one-dimensional marginal posterior of the luminosity distance. A distribution is classified as unimodal if the $p$-value exceeds 0.01. As shown in Figure 10, about 19% (23%) of sources have a multimodal $D_{\mathrm{L}}$ posterior for 2L MisA (2L A), while the $\Delta$ configurations display only $\sim 2$%.
Appendix G Sky modes in the 2L Misaligned configuration
Figure 11: Marginalized one- and two-dimensional posterior distributions for all parameters, with the exception that this time we show the source-frame chirp mass ($\mathcal{M}_{\mathrm{s}}$) and redshift ($z$). The results for the low-mass event (orange) are shown alongside those for the high-mass event (blue). Vertical and horizontal lines indicate the true injected values and contours represent the 68% and 95% credible regions. Top-right panel: Sky map of the low-mass (orange) and high-mass (blue) events, with contours indicating the 50% and 90% credible area. See Appendix G for further details.
Figure 2 illustrates multimodalities in sky localization for a source observed with the 2L MisA configuration. The presence of these multimodalities is further confirmed in Figure 4, where almost no sources exhibit a single mode. Most sources ($\gtrsim 60\%$) display between two and six sky modes, while $\sim 20\%$ exhibit eight distinct modes.
In this section, we provide an intuitive example to illustrate that the presence of multiple sky modes in the 2L MisA configuration is specific to the sources considered in this study, after applying the sample efficiency selection (Section III.1 and bottom panel of Figure 1). For a small subset of sources that are excluded from our analysis but still present in the prior space, it is reasonable to expect that the number of sky modes would be significantly lower. Intuitively, the baseline between the 2L ET interferometers becomes effectively long for sources with sufficiently high SNR above a certain frequency threshold, which depends on both the length of baseline and the source’s sky location. A comprehensive study to quantify this effect is beyond the scope of this manuscript and will be addressed in future work.
Figure 11 shows the one- and two-dimensional marginal distributions obtained with Bilby for two sources at zero noise: a less massive, closer source with $\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}=55.5~\mathrm{M}_{\odot}$ and $D^{\mathrm{inj}}_{\mathrm{L}}=6.2~\mathrm{Gpc}$, with optimal SNR = 87.5, and a more massive, more distant source with $\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}=145.3~\mathrm{M}_{\odot}$ and $D^{\mathrm{inj}}_{\mathrm{L}}=37.2~\mathrm{Gpc}$ with optimal SNR = 28.7. The sky location, source-frame chirp mass, mass ratio, and other parameters are fixed for both sources to the maximum-likelihood values estimated for GW250114 (LIGOScientific:2025rid), while only the luminosity distance—and thus the redshift—are varied.
Figure 11 illustrates that multimodalities in sky localization vanish for the less massive source, whereas they persist for the more massive one, yielding $\Delta\Omega_{90\%}=3$ deg2 for the low-mass event and $\Delta\Omega_{90\%}=194$ deg2 for the high-mass event.
Appendix H Performance in estimating other parameters
Figure 12: Cumulative distributions of events as a function of the optimal SNR ($\rho$, Equation D4), the relative variations in detector-frame chirp mass ($\Delta\mathcal{M}_{\mathrm{d}}/\mathcal{M}^{\mathrm{inj}}_{\mathrm{d}}$), mass ratio ($q$), first ($\Delta\chi_{1}$) and second ($\Delta\chi_{2}$) aligned spins, and inclination angle ($\Delta\theta_{\mathrm{jn}}$) for all considered XG detector configurations. See Appendix H for details.
For completeness, Figure 12 presents the performance of the considered detector configurations in estimating the remaining parameters, such as optimal SNR, detector-frame chirp mass, mass ratio, the aligned spin components, and inclination angle. The overall trend is confirmed, as the $\Delta$ and 2L A configurations provide the poorest performance, while the 2L MisA + CE setup delivers the most accurate estimates..